# WorldGUI: An Interactive Benchmark for Desktop GUI Automation from Any Starting Point

### v3

@@ Abstract
GUI agents have achieved outstanding performance in GUI element grounding. However, planning remains highly challenging, especially due to the sensitivity to the initial state of the environment. Specifically, slight differences in the initial state-such as the target software not being open or the interface not being in its default state, often lead to planning errors. This issue is widespread in real application scenarios, but existing benchmarks fail to evaluate it. To address this gap, we introduce WorldGUI, a comprehensive GUI benchmark containing tasks across ten widely used desktop and web applications (e.g., PowerPoint, VSCode, Acrobat), each instantiated with diverse initial states to simulate authentic human–computer interactions. Complementing this, we propose WorldGUI-Agent, a universal framework that unifies three core modules: Planner-Critic for high-level plan refinement, Step-Check for intermediate verification, and Actor-Critic for action-level optimization to proactively detect and correct errors. Experimental evaluation shows that WorldGUI-Agent outperforms the outstanding existing model (Claude-3.5 Computer Use) by 12.4% in success rate on WorldGUI, and achieves a 31.2% overall success rate on WindowsAgentArena, surpassing the prior state-of-the-art by 11.7%. Our analysis further reveals that dynamic augmentation tasks and desktop environments pose substantial hurdles, underscoring the necessity of adaptive planning and feedback-driven execution for advancing real-world GUI automation. The code and data are available at https://github.com/showlab/WorldGUI.
00footnotetext: Corresponding author.
Figure 1: Software taxonomy of WorldGUI and the performance comparison of GUI agents. The left shows 5 main groups and 10 software in our WorldGUI. The right shows that WorldGUI-Agent surpasses previous SOTA GUI agents significantly.
@@ 1 Introduction
Graphical User Interface (GUI) automation has become a prominent research area, driven by the need to enhance user productivity. This domain encompasses software usage, file management, office design, coding, and web browsing. Building upon Multimodal Large Language Models (MLLMs) such as GPT-4o gpt4o and Claude-3.5 claude3.5 , GUI agents have the potential to solve various computer tasks to avoid repetitive work or as an AI assistant to enhance productivity efficiency.
GUI automation operates in a dynamic environment, which goes beyond the traditional computer vision tasks like image recognition he2016deep and visual question answering antol2015vqa ; VQAv2 . However, current online GUI benchmarks such as WebArena webarena , WebVoyager webvoyager , and WindowsAgentArena windowsagentarena do not capture this dynamism. Currently, most GUI benchmarks OSWorld ; windowsagentarena ; assistgui ; webarena ; visualwebarena ; webvoyager focus on initial and final states, measuring success rates but overlooking the state variety in real GUI scenarios. These benchmarks often ignore situations where: (1) The software interface is not in its default state. (2) The human-computer interactions may start from the intermediate state of a specific task. (3) Differences in agent robustness, where agents with the same low success rate (e.g., 20%) may vary in their ability to self-reflection, but these abilities cannot be measured in a static setting. As a result, these benchmarks fail to comprehensively assess the GUI agents.
Figure 2: WorldGUI. The left shows that for each task, WorldGUI provides a user query, instructional video, and pre-actions. The pre-actions lead to different initial states. The key characteristic of our WorldGUI is the various initial states of the same task to stimulate the real-world testing process. The right shows the software included in our benchmark and the interactions about testing the agents in our GUI environment.
In this paper, we take the first step toward comprehensive GUI evaluation by designing GUI tasks with various initial states. We consider that the testing process of WorldGUI can be featured: (1) Intermediate Starting States: Real user interactions with GUI assistants do not always begin from default initial conditions, allowing tasks to start from intermediate states where users may seek assistance at any point. (2) Contextual Variability: In some cases, tasks may originate from entirely different contexts or interfaces, requiring the agent to adapt by modifying existing plans or introducing new steps to ensure task execution. By incorporating these situations into the benchmark design, WorldGUI better mirrors real-world GUI interactions, enabling a more accurate and thorough assessment of GUI agent capabilities. Specifically, WorldGUI embraces 10 widely-used desktop applications with 611 tasks in total. For each task, we create a user query, an instructional video, and the corresponding project file. We engaged four trained annotators skilled in using these applications for annotation. To stimulate the dynamic testing scenarios, we demonstrate each task to obtain ground-truth (GT) plans and then conduct the augmentations for each task using pre-actions.
In addition, we introduce a new GUI agent framework, WorldGUI-Agent, which builds upon critical thinking design principle, an aspect less emphasized in previous GUI agents hong2024cogagent ; cheng2024seeclick ; autowebglm ; AgentS ; osatlas . In dynamic GUI environments, application settings may not be in default configurations. This unpredictability requires agents to have the essential ability to detect and adapt to such changes to ensure task accuracy. Through our analysis of real-world GUI scenarios, we identify three design principles for GUI agents: (1) Post-Planning Critique, (2) Pre-Action Validation, and (3) Post-Action Evaluation. We argue that these components are fundamental and universal for GUI agents.
To summarize, our key contributions are the following: (1) We are the first to stress the dynamic testing processes in the online GUI testing and propose a new benchmark called WorldGUI; (2) We introduce WorldGUI-Agent, a fundamental and universal GUI framework that incorporates critical thinking into the overall agent design, providing valuable insight and guidance for future development; (3) We explore the essential property of critical thinking in GUI agents and empirically show that critical thinking is extremely useful for handling GUI tasks (see Figure 1).

### v4

@@ Abstract
Recent progress in GUI agents has substantially improved visual grounding, yet robust planning remains challenging, particularly when the environment deviates from a canonical initial state. In real applications, users often invoke assistance mid-workflow, where software may be partially configured, steps may have been executed in different orders, or the interface may differ from its default setup. Such task-state variability is pervasive but insufficiently evaluated in existing GUI benchmarks. To address this gap, we introduce WorldGUI, a benchmark covering ten widely used desktop and web applications with tasks instantiated under diverse, systematically constructed initial states. These variations capture realistic human–computer interaction settings and enable diagnostic evaluation of an agent’s ability to recover, adapt plans, and handle non-default contexts. We further present WorldGUI-Agent, a simple and model-agnostic framework that organizes planning and execution around three critique stages, improving reliability in dynamic environments. Experiments demonstrate that state-of-the-art GUI agents exhibit substantial performance degradation under non-default initial conditions, revealing limited robustness and fragile planning behaviors. Our benchmark and framework provide a foundation for developing more adaptable and reliable GUI agents. The code and data are available at https://github.com/showlab/WorldGUI.
@@ 1 Introduction
Table 1: Comparison with other online GUI benchmarks. WorldGUI is a unique benchmark that embraces diverse initial states and better reflects the authentic interactions in GUI scenarios. Env?: Indicates whether an environment is required to be deployed.
Benchmark Software Task Platform Env? Inst. Video? GT Plan Task State
variability?
WebArena (Zhou et al., ) 6 812 Web Yes
VisualWebArena (Koh et al., 2024) 3 910 Web Yes
WebVoyager (He et al., 2024) 15 643 Web Yes
AutoDroid (Wen et al., 2024) 13 158 Android OS Yes
AndroidWorld (Rawles et al., 2024) 20 116 Android OS Yes
AgentStudio (Zheng et al., 2025) 9 205 Desktop + Web Yes
Mobile-Eval (Wang et al., 2024) 10 30 Android OS Yes
APPAgent (Zhang et al., 2023) 10 50 Android OS Yes
OSWorld (Xie et al., 2024) 10 369 Desktop Yes
AssistGUI (Gao et al., 2024) 9 100 Windows No
WindowAgentArena (Bonatti et al., 2024) 11 154 Windows Yes
WorldGUI 10 611 Win. + Web No
2 Introduction
@@
Graphical User Interface (GUI) automation has emerged as a promising direction for productivity-oriented AI systems. With the rapid advancement of Multimodal Large Language Models (MLLMs) such as GPT-5 (Singh et al., 2025) and Claude-4-Sonnet (Anthropic, 2025), GUI agents have demonstrated the potential to assist users in software usage, file management, office design, coding, and web browsing, reducing repetitive work and improving workflow efficiency.
Unlike traditional computer vision tasks such as image recognition (He et al., 2016) or visual question answering (Antol et al., 2015; Goyal et al., 2017), GUI automation operates in a highly dynamic environment where the interface state evolves as a result of user actions, system configurations, or partial task progress. However, existing online GUI benchmarks, including WebArena (Zhou et al., ), WebVoyager (He et al., 2024), OSWorld (Xie et al., 2024), AssistGUI (Gao et al., 2024), and WindowsAgentArena (Bonatti et al., 2024), primarily evaluate agents from a single, canonical initial state and measure success only at the end of the trajectory. Such static evaluation overlooks several common situations in real-world usage: (1) software interfaces are often not in their default configurations, (2) users frequently request assistance from intermediate states of partially completed tasks, and (3) agents with the same low success rate (e.g., 20%) may differ substantially in their robustness, planning stability, and self-correction behavior differences that cannot be revealed under a single start-state protocol.
In parallel, robustness-oriented GUI benchmarks such as GUI-Robust (Yang et al., 2025) examine how agents respond to interface-level perturbations, including layout shifts, appearance variations, and transient pop-ups. These benchmarks vary the visual presentation while keeping the task goal and initial state fixed, providing valuable insights into perception robustness. In contrast, WorldGUI does not alter UI skin, resolution, or layout; instead, it targets task-progress–level variation by modifying the executed steps themselves (Add-, Trim-, and Adjust-step). This form of task-state variability, such as missing, redundant, or reordered steps, poses challenges fundamentally different from interface perturbations, making the two robustness dimensions complementary and largely orthogonal.
To complement existing benchmarks, we focus on robustness to intermediate starting states, a ubiquitous but under-explored aspect of GUI interactions. We introduce WorldGUI, a benchmark that systematically constructs diverse initial states for each task by applying controlled pre-actions. These pre-actions modify task progress itself (e.g., steps partially completed, over-executed, or executed in a different order), allowing tasks with the same goal to begin from multiple realistic contexts. This design mirrors how real users summon assistance in the middle of workflows and enables diagnostic evaluation of an agent’s ability to recover, adapt plans, and handle contextual variability. WorldGUI covers 10 widely used desktop applications and provides 611 tasks, each paired with a user query, instructional video, and project file. To construct meaningful state variations, trained annotators first demonstrate ground-truth (GT) plans, after which we generate augmented initial states via pre-actions.
In addition, we propose WorldGUI-Agent, a simple and general GUI agent framework grounded in critical-thinking principles, an aspect that has received relatively limited attention in prior GUI agents (Hong et al., 2024; Cheng et al., 2024; Lai et al., 2024; Agashe et al., 2025a; Wu et al., 2024). Dynamic GUI environments often deviate from expected states due to user behavior or system configurations, making it essential for agents to detect mismatches and adapt their plans. Through an analysis of real-world GUI scenarios, we distill three design principles that we argue are fundamental for GUI automation: (1) Post-Planning Critique, (2) Pre-Action Validation, and (3) Post-Action Evaluation. These components enable WorldGUI-Agent to refine plans, verify step prerequisites, and evaluate action outcomes, thereby improving reliability in dynamic settings.
Experiments reveal three key observations: (1) a substantial gap remains between humans and current agents, with models such as UI-Tars and GPT-5.1 still struggling on our curated desktop GUI tasks; (2) performance degrades sharply under non-default initial states, highlighting limited robustness and fragile planning capabilities; and (3) WorldGUI-Agent significantly improves robustness and overall task success across WorldGUI and WindowsAgentArena, even when built upon weaker base models and without prompt tuning, demonstrating the importance of multi-level critique for dynamic GUI automation (see Section 6).
3 Related Work
Figure 1: WorldGUI. Left: WorldGUI creates pre-actions for each meta task, leading to different initial states. It successfully reflects the real-world human-computer interaction process. Right: components in WorldGUI.
3.1 GUI Benchmarks
GUI benchmarks provide the foundation for evaluating computer-use agents across web, desktop, and mobile environments. Early interactive benchmarks such as WebShop (Yao et al., 2022), WebArena (Zhou et al., ), and WebVoyager (He et al., 2024) center on browser tasks with long-horizon goals. OSWorld (Xie et al., 2024) and mobile benchmarks, including MobileAgent (Wang et al., 2024) and AppAgent (Zhang et al., 2023) broaden coverage to multi-application desktop and mobile ecosystems. AssistGUI (Gao et al., 2024) and WindowsAgentArena (Bonatti et al., 2024) provide online execution environments for Windows-based tasks.
Recent benchmarks emphasize scalability and online evaluation. UIExplorer (Nica et al., 2025) offers a large-scale environment for GUI exploration in both structured-DOM and pure screen modes. GUI-World (Chen et al., 2025) provides multi-application desktop tasks with online execution and standardized initial states. Robustness-oriented suites such as GUI-Robust (Yang et al., 2025) examine agents under interface-level perturbations (e.g., layout shifts, visual noise, pop-ups), focusing on perception and action robustness. While these benchmarks significantly advance GUI automation, they overwhelmingly assume a single canonical initial state for each task. They do not explicitly evaluate task-state variability, such as partially completed workflows, missing or redundant steps, or alternative intermediate contexts common in real user workflows. Our benchmark, WorldGUI, complements existing online environments by systematically constructing diverse initial states for the same task goal via controlled pre-actions, enabling diagnostic evaluation of robustness, recoverability, and planning stability in dynamic GUI environments.
3.2 GUI Agents
Recent progress in GUI agents spans perception, grounding, planning, and full computer-use control. CogAgent (Hong et al., 2024) improves GUI perception, while SeeClick (Cheng et al., 2024) and SeeAct (Zheng et al., 2024) focus on grounding UI elements for more reliable action prediction. MobileAgent (Wang et al., 2024) and AppAgent (Zhang et al., 2023) extend these capabilities to mobile environments. On desktop and OS-level tasks, AssistGUI (Gao et al., 2024) adopts a Plan-Act pipeline with a single post-action checker. Agent-S (Agashe et al., 2025a) introduces experience-augmented hierarchical planning with a Manager-Worker architecture, a self-evaluator, and an Agent-Computer Interface (ACI). Agent-S2 (Agashe et al., 2025b) further incorporates specialist grounding (Mixture-of-Grounding) and Proactive Hierarchical Planning, achieving strong performance across OSWorld, WindowsAgentArena, and AndroidWorld. Beyond inference-time frameworks, several works train GUI agents or their components. Aguvis (Xu et al., 2024) and UGround (Gou et al., 2025) train pure-vision or universal grounding models at scale. AgentTrek (Xu et al., 2025) synthesizes multi-step trajectories from web tutorials for supervised training, and SE-GUI (Yuan et al., 2025) applies reinforcement learning to strengthen grounding robustness. While these approaches enhance perception, grounding, or hierarchical planning, they typically employ only a single verification stage (e.g., self-evaluation or post-action checking). In contrast, our framework WorldGUI-Agent focuses on structuring the verification process itself for dynamic GUI environments. WorldGUI-Agent introduces three complementary critique stages: Post-Planning Critique, Pre-Action Validation, and Post-Action Evaluation, enabling agents to detect state mismatches, assess step readiness, and repair execution failures. This multi-level critique design is orthogonal to hierarchical planners and training-based pipelines, and can be combined with them without additional model training. A component-wise comparison is provided in Table 3.